OpenStudy (anonymous):

Which polygon has an interior measure of 900 degrees? A hexagon B heptagon C octagon D nonagon

OpenStudy (johnweldon1993):

Hint... \[\large (n - 2) \times 180\] this tells us for a polygon of 'n' sides...the total measure of the interior angles... so if we need 900 \[\large (n - 2) \times 180 = 900\] \[\large n - 2 = 5\] \[\large n = 7\] Also known as a Heptagon!